# Kernel Ridge Regression with Predicted Feature Inputs and Applications to Factor-Based Nonparametric Regression

### v1

@@ 5 Extension to other loss functions
Appendix A contains numerical studies of the kernel ridge regression with features predicted by the principal component analysis under models (3) and (4). The main proofs are collected in Appendix B and presented on a section-by-section basis. Auxiliary lemmas are stated in Appendix C.
Appendix A Numerical experiments
In this section, we evaluate the numerical performance of the proposed algorithm in Section 2 for KRR with predicted feature inputs, and compare it with several competitors. Specifically, we consider the following nonparametric latent factor model:
$Y=f^{*}(Z)+\epsilon,$ (58)
$X=AZ+W.$ (59)
where the latent vector $Z=(Z_{1},Z_{2},Z_{3})^{\top}$, with $r=3$, is generated with entries independently sampled from $\text{Unif}(0,1)$, the true regression function $f^{*}$ is set as
$f^{*}(Z)=2\sin\big{(}3\pi Z_{1}\big{)}+3|Z_{1}-0.5|-\exp\big{(}Z_{2}^{2}-Z_{3}^{2}\big{)},$
the regression error $\epsilon\sim N(0,0.8^{2})$, the loading matrix $A$ is generated with rows independently sampled from $N(0,{\rm diag}(10,5.5,1))$ and $W$ is generated with entries independently sampled from $N(0,1.5^{2})$.
In each setting, we generate the training data $\mathcal{T}=\{(Y_{i},X_{i},Z_{i})\}_{i=1}^{n}$ i.i.d. according to models (58) and (59). We also generate another auxiliary data $\mathbf{X}^{\prime}=(X^{\prime}_{1},\ldots,X^{\prime}_{n})^{\top}$ independently from model (59). Under model (59), as illustrated in Section 4, the latent factors can be predicted by
where $\widehat{B}^{\prime}$ is constructed from $\mathbf{X}^{\prime}$ via PCA as in (50). According to our theory, the auxiliary data $\mathbf{X}^{\prime}$ is needed to construct $\widehat{B}$ that is independent of $\mathcal{T}$. Numerically, instead of using $\mathbf{X}^{\prime}$, we can also predict $\mathbf{Z}$ by $\widehat{\mathbf{Z}}=(\widehat{B}X_{1},\ldots,\widehat{B}X_{n})^{\top}$, where $\widehat{B}$ is obtained by applying PCA on $\mathbf{X}=(X_{1},\ldots,X_{n})^{\top}$. For comparison, we consider the following predictors:
  • KRR-$\widehat{\mathbf{Z}}^{\prime}$: KRR in (12) with predicted inputs using $\widehat{\mathbf{Z}}^{\prime}$;
  • KRR-$\widehat{\mathbf{Z}}$: KRR in (12) with predicted inputs using $\widehat{\mathbf{Z}}$;
  • KRR-$\mathbf{Z}$: KRR in (12) using the true $\mathbf{Z}$;
  • KRR-$\mathbf{X}$: Classic KRR obtained from (12) with $X_{i}$ in place of $\widehat{g}(X_{i})$;
  • LR-$\mathbf{Z}$: Regressing $\mathbf{Y}$ onto $\mathbf{Z}$ by OLS.
KRR-$\mathbf{X}$ is included for comparing the proposed KRR-$\widehat{\mathbf{Z}}^{\prime}$ with the classic KRR method. The linear predictor LR-$\mathbf{Z}$ is considered to illustrate the benefit of using non-linear predictors. KRR-$\mathbf{Z}$ uses the unobserved latent factors $\mathbf{Z}$ to illustrate the additional prediction errors of KRR-$\widehat{\mathbf{Z}}^{\prime}$ due to predicting $\mathbf{Z}$. Finally, KRR-$\widehat{\mathbf{Z}}$ is included to examine the effect of using an auxiliary data set to compute $\widehat{B}$.
For the kernel-based methods, the regularization parameter $\lambda$ in (12) is chosen via 3-fold cross-validation and we choose the RKHS induced by the Gaussian kernel, $K(z,z^{\prime})=\exp(-\|z-z^{\prime}\|_{2}^{2}/\sigma^{2}_{n})$, where $\sigma_{n}$ is set as the median of all pairwise distances among sample points, as suggested in Mukherjee et al. (2010).
To evaluate the numerical performance, a test dataset $\mathcal{T}^{t}=\{(Y_{i}^{t},X_{i}^{t},Z_{i}^{t})\}_{i=1}^{m}$ is also generated independently from models (58) and (59). For any predictor $\widehat{f}:\mathbb{R}^{p}\to\mathbb{R}$, we evaluate its predictive performance by
$\mathbb{E}_{m}\big{[}Y-\widehat{f}(X)\big{]}^{2}:=\frac{1}{m}\sum_{i=1}^{m}\left(Y^{t}_{i}-\widehat{f}(X^{t}_{i})\right)^{2}.$
For any predictor $\widehat{f}:\mathbb{R}^{r}\to\mathbb{R}$ using the true latent factor $\mathbf{Z}$, we evaluate its performance by
$\mathbb{E}_{m}\big{[}Y-\widehat{f}(Z)\big{]}^{2}:=\frac{1}{m}\sum_{i=1}^{m}\left(Y^{t}_{i}-\widehat{f}(Z^{t}_{i})\right)^{2}.$
These metrics are computed based on $100$ repetitions in each setting.
A.1 The effect of training sample size
We first study how the predictive performance of different predictors changes when the number of training samples varies as $n\in\{200,600,1000,1400,1800,2200\}$. For each $n$, we fix $p=2000$ and $n=m$. The averaged empirical prediction errors, along with the corresponding standard deviations (included in parentheses), are reported in Table 2.
Table 2: The averaged performance of all the predictors with varying $n$.
Sample size $n$ 200 600 1000 1400 1800 2200
KRR-$\widehat{\mathbf{Z}}^{\prime}$ 1.59 0.90 0.82 0.79 0.77 0.75
(0.32) (0.06) (0.04) (0.04) (0.03) (0.02)
KRR-$\widehat{\mathbf{Z}}$ 1.63 0.90 0.81 0.79 0.77 0.75
(0.32) (0.07) (0.04) (0.03) (0.03) (0.02)
KRR-$\mathbf{Z}$ 1.08 0.77 0.72 0.71 0.70 0.68
(0.19) (0.05) (0.03) (0.03) (0.03) (0.02)
KRR-$\mathbf{X}$ 1.99 1.91 1.89 1.85 1.83 1.82
(0.19) (0.10) (0.08) (0.07) (0.06) (0.05)
LR-$\mathbf{Z}$ 3.61 3.58 3.62 3.60 3.61 3.60
(0.29) (0.16) (0.13) (0.10) (0.11) (0.08)
As shown in Table 2, the proposed KRR-$\widehat{\mathbf{Z}}^{\prime}$ outperforms all other predictors, except for the oracle approach, KRR-$\mathbf{Z}$, and its variant KRR-$\widehat{\mathbf{Z}}$ that does not use the auxiliary data. KRR-$\widehat{\mathbf{Z}}^{\prime}$ has increasingly closer performance to KRR-$\mathbf{Z}$ as the sample size grows. This is in line with our risk bound in Corollary 5 where the extra prediction error of KRR-$\widehat{\mathbf{Z}}^{\prime}$ relative to KRR-$\mathbf{Z}$ is quantified by $\mathcal{O}_{\mathbb{P}}(1/p+1/n)$, which reduces to $1/p$ as $n$ increases. Comparing to KRR-$\widehat{\mathbf{Z}}$, using auxiliary data only seems to have benefit for small sample sizes. Compared to KRR-$\widehat{\mathbf{Z}}^{\prime}$ and KRR-$\mathbf{Z}$, increasing the sample size has a much smaller effect in reducing the prediction error of KRR-$\mathbf{X}$, indicating a slower rate of convergence of the latter on the sample size. This is expected, as KRR-$\mathbf{X}$ does not exploit the low-dimensional structure. Finally, comparing to the predictor based on OLS in the second step, KRR-$\widehat{\mathbf{Z}}^{\prime}$ has substantial predictive advantage, implying the benefit of using KRR to capture the non-linear relationship between the response and the latent factor.
A.2 The effect of dimensionality
We next examine the predictors when the number of features varies as $p\in\{50,100,500,1000,1500,3000\}$. For each $p$, we set $n=m=800$. As shown in Table 3, the predictive performance between KRR-$\widehat{\mathbf{Z}}^{\prime}$ and its variant, KRR-$\widehat{\mathbf{Z}}$, are nearly identical. KRR-$\widehat{\mathbf{Z}}^{\prime}$ has also the close performance to KRR-$\mathbf{Z}$ and their difference gets smaller as $p$ increases. Indeed, as shown in Corollary 5, the prediction error of KRR-$\widehat{\mathbf{Z}}^{\prime}$, due to predicting $\mathbf{Z}$, contains the term $1/p$ which decreases as $p$ grows. Since KRR-$\mathbf{Z}$ and LR-$\mathbf{Z}$ use $\mathbf{Z}$ instead of $\mathbf{X}$, their prediction errors remain the same as $p$ increases. We also see that the performance of KRR-$\mathbf{X}$ slightly increases as $p$ grows. One explanation could be that more useful information for predicting $Y$ can be found from $\mathbf{X}$ as $p$ increases. The linear predictor that uses PCs still suffers from underfitting.
Table 3: The averaged performance of all the predictors with varying $p$.
Dimension $p$ 50 100 500 1000 1500 3000
KRR-$\widehat{\mathbf{Z}}^{\prime}$ 1.85 1.44 0.95 0.89 0.87 0.84
(0.15) (0.12) (0.06) (0.06) (0.05) (0.05)
KRR-$\widehat{\mathbf{Z}}$ 1.86 1.43 0.94 0.89 0.87 0.83
(0.16) (0.12) (0.06) (0.06) (0.05) (0.05)
KRR-$\mathbf{Z}$ 0.75 0.74 0.74 0.75 0.75 0.74
(0.04) (0.04) (0.04) (0.03) (0.04) (0.04)
KRR-$\mathbf{X}$ 2.32 2.14 1.95 1.94 1.90 1.83
(0.17) (0.13) (0.10) (0.12) (0.10) (0.09)
LR-$\mathbf{Z}$ 3.60 3.60 3.59 3.62 3.60 3.60
(0.15) (0.14) (0.14) (0.12) (0.15) (0.14)
A.3 The effect of signal-to-noise ratio
We investigate how the signal-to-noise ratio (SNR) of predicting the latent factor $Z$, defined as
$\text{SNR}:=\lambda_{r}(A\Sigma_{Z}A^{\top})/\|\Sigma_{W}\|_{\rm op},$
affects the prediction performance of different predictors. A higher SNR implies more information of $\mathbf{Z}$ contained in $\mathbf{X}$, which leads to better prediction of $\mathbf{Z}$ and improves the prediction of $Y$. So it is expected that only KRR-$\widehat{\mathbf{Z}}^{\prime}$, KRR-$\widehat{\mathbf{Z}}$ and KRR-$\mathbf{X}$ behave differently as the SNR changes. In this experiment, we vary the SNR by multiplying $A$ by a scalar $\alpha$ within $\{0.1,0.2,0.3,0.5,0.75,1,1.5\}$. In all settings, we fix $n=m=800$ and $p=2000$. The obtained results are reported in Table 4.
As expected, increasing SNR leads to better prediction performance for KRR-$\widehat{\mathbf{Z}}^{\prime}$, KRR-$\widehat{\mathbf{Z}}$ and KRR-$\mathbf{X}$ while the prediction errors of other predictors remain unchanged. However, KRR-$\mathbf{X}$ continues to perform poorly even for the settings with large values of SNR due to underfitting. On the contrary, when the SNR is sufficiently large, the prediction error of KRR-$\widehat{\mathbf{Z}}^{\prime}$ becomes very small, approaching that of KRR-$\mathbf{Z}$. We also observe from the comparison between KRR-$\widehat{\mathbf{Z}}^{\prime}$ and KRR-$\widehat{\mathbf{Z}}$ that the benefit of using auxiliary data becomes more visible for small SNRs.
Table 4: The averaged performance of all the predictors with varying $\alpha$.
SNR parameter $\alpha$ 0.1 0.2 0.3 0.5 0.75 1 1.5
KRR-$\widehat{\mathbf{Z}}^{\prime}$ 2.59 1.46 1.16 0.94 0.87 0.85 0.82
(0.13) (0.10) (0.07) (0.06) (0.06) (0.05) (0.05)
KRR-$\widehat{\mathbf{Z}}$ 2.83 1.56 1.14 0.93 0.86 0.84 0.82
(0.15) (0.13) (0.06) (0.06) (0.05) (0.05) (0.05)
KRR-$\mathbf{Z}$ 0.74 0.75 0.74 0.74 0.74 0.74 0.74
(0.04) (0.04) (0.04) (0.04) (0.04) (0.04) (0.04)
KRR-$\mathbf{X}$ 3.79 3.74 3.47 2.53 2.11 1.88 1.39
(0.16) (0.16) (0.14) (0.11) (0.10) (0.08) (0.07)
LR-$\mathbf{Z}$ 3.62 3.59 3.61 3.59 3.58 3.61 3.61
(0.14) (0.15) (0.14) (0.14) (0.13) (0.15) (0.15)
Appendix B Main Proofs

### v2

@@ 5 Extension to other loss functions
6 Simulations
In this section, we evaluate the predictive performance of KRR using predicted features. We consider models (3) and (4) where the latent vector $Z=(Z_{1},Z_{2},Z_{3})^{\top}$, with $r=3$, is generated with entries independently sampled from $\text{Unif}(0,1)$, the regression function $f^{*}$ is set as
$f^{*}(Z)=2\sin(3\pi Z_{1})+3|Z_{1}-0.5|-\exp(Z_{2}^{2}-Z_{3}^{2}),$
the regression error $\epsilon\sim N(0,0.8^{2})$, the loading matrix $A$ is generated with rows independently sampled from $N(0,{\rm diag}(10,5.5,1))$ and $W$ is generated with entries independently sampled from $N(0,1.5^{2})$. The training data $\mathcal{T}=\{(Y_{i},X_{i},Z_{i})\}_{i=1}^{n}$, the auxiliary data $\mathbf{X}^{\prime}=(X^{\prime}_{1},\ldots,X^{\prime}_{n})^{\top}$ and the test dataset $\mathcal{T}^{t}=\{(Y_{i}^{t},X_{i}^{t},Z_{i}^{t})\}_{i=1}^{m}$ are generated i.i.d. according to models (3) and (4).
Under model (3), as discussed in Section 4, we predict $\mathbf{Z}=(Z_{1},\ldots,Z_{n})^{\top}$ by
with $\widehat{B}^{\prime}$ constructed from $\mathbf{X}^{\prime}$ via PCA as in (45). We also consider predicting $\mathbf{Z}$ by
$\widehat{\mathbf{Z}}=(\widehat{B}X_{1},\ldots,\widehat{B}X_{n})^{\top}$
with $\widehat{B}$ obtained by applying PCA on $\mathbf{X}=(X_{1},\ldots,X_{n})^{\top}$. The following predictors are considered in comparison: (1) $\widehat{\mathbf{Z}}^{\prime}: KRR$ in (12) with predicted inputs using $\widehat{\mathbf{Z}}^{\prime}$; (2) $\widehat{\mathbf{Z}}: KRR$ in (12) with predicted inputs using $\widehat{\mathbf{Z}}$; (3) $\mathbf{Z}: KRR$ in (12) using the true $\mathbf{Z}$; (4) $\mathbf{X}: KRR$ from (12) with $X_{i}$ in place of $\widehat{g}(X_{i})$; and (5) LR-$\mathbf{Z}$: Regressing $\mathbf{Y}$ onto $\mathbf{Z}$ by OLS. Note that KRR-$\mathbf{X}$ is included for comparing the proposed KRR-$\widehat{\mathbf{Z}}^{\prime}$ with the classical KRR method. The linear predictor LR-$\mathbf{Z}$ is considered to illustrate the benefit of using non-linear predictors. KRR-$\mathbf{Z}$ uses the true $\mathbf{Z}$ to illustrate the additional prediction errors of KRR-$\widehat{\mathbf{Z}}^{\prime}$ due to predicting $\mathbf{Z}$. Finally, KRR-$\widehat{\mathbf{Z}}$ is included to examine the effect of using an auxiliary data set to compute $\widehat{B}$.
For the kernel-based methods, we select the regularization parameter $\lambda$ in (12) via 3-fold cross-validation and choose the Gaussian kernel, $K(z,z^{\prime})=\exp(-\|z-z^{\prime}\|_{2}^{2}/\sigma^{2}_{n})$, with $\sigma_{n}$ set as the median of all pairwise distances among sample points, as suggested in Mukherjee et al. (2010). For any predictor $\widehat{f}:\mathbb{R}^{p}\to\mathbb{R}$, we evaluate its predictive performance by $m^{-1}\sum_{i=1}^{m}(Y^{t}_{i}-\widehat{f}(X^{t}_{i}))^{2}.$ For any predictor $\widehat{f}:\mathbb{R}^{r}\to\mathbb{R}$ using the true $\mathbf{Z}$, we evaluate its performance by $m^{-1}\sum_{i=1}^{m}(Y^{t}_{i}-\widehat{f}(Z^{t}_{i}))^{2}$. These metrics are computed based on $100$ repetitions in each setting.
We examine the effect of the following parameters on the predictive performance by varying them one at a time: (1) the sample size $n\in\{200,600,1000,1400,1800,2200\}$; (2) the number of features $p\in\{50,100,500,1000,1500,3000\}$; (3) the signal-to-noise ratio (SNR) of predicting the latent factor $Z$ in the presence of $W$, defined as $\text{SNR}:=\lambda_{r}(A\Sigma_{Z}A^{\top})/\|\Sigma_{W}\|_{\rm op}$, by multiplying $A$ by a scalar $\alpha\in\{0.1,0.2,0.3,0.5,0.75,1,1.5\}$. We fix $n=800$, $p=2000$ and $\text{SNR}=1$ whenever they are not varied, and set $m=n$ in all settings. The averaged prediction errors in each setting are reported in Table 3.
Table 3: The averaged performance of all the predictors with varying $n$, $p$, and $\alpha$.
Sample size $n$ 200 600 1000 1400 1800 2200
KRR-$\widehat{\mathbf{Z}}^{\prime}$ 1.59 (0.32) 0.90 (0.06) 0.82 (0.04) 0.79 (0.04) 0.77 (0.03) 0.75 (0.02)
KRR-$\widehat{\mathbf{Z}}$ 1.63 (0.32) 0.90 (0.07) 0.81 (0.04) 0.79 (0.03) 0.77 (0.03) 0.75 (0.02)
KRR-$\mathbf{Z}$ 1.08 (0.19) 0.77 (0.05) 0.72 (0.03) 0.71 (0.03) 0.70 (0.03) 0.68 (0.02)
KRR-$\mathbf{X}$ 1.99 (0.19) 1.91 (0.10) 1.89 (0.08) 1.85 (0.07) 1.83 (0.06) 1.82 (0.05)
LR-$\mathbf{Z}$ 3.61 (0.29) 3.58 (0.16) 3.62 (0.13) 3.60 (0.10) 3.61 (0.11) 3.60 (0.08)
Dimension pp 50 100 500 1000 1500 3000
KRR-$\widehat{\mathbf{Z}}^{\prime}$ 1.85 (0.15) 1.44 (0.12) 0.95 (0.06) 0.89 (0.06) 0.87 (0.05) 0.84 (0.05)
KRR-$\widehat{\mathbf{Z}}$ 1.86 (0.16) 1.43 (0.12) 0.94 (0.06) 0.89 (0.06) 0.87 (0.05) 0.83 (0.05)
KRR-$\mathbf{Z}$ 0.75 (0.04) 0.74 (0.04) 0.74 (0.04) 0.75 (0.03) 0.75 (0.04) 0.74 (0.04)
KRR-$\mathbf{X}$ 2.32 (0.17) 2.14 (0.13) 1.95 (0.10) 1.94 (0.12) 1.90 (0.10) 1.83 (0.09)
LR-$\mathbf{Z}$ 3.60 (0.15) 3.60 (0.14) 3.59 (0.14) 3.62 (0.12) 3.60 (0.15) 3.60 (0.14)
SNR $\alpha$ 0.1 0.2 0.3 0.5 0.75 1 1.5
KRR-$\widehat{\mathbf{Z}}^{\prime}$ 2.59 (0.13) 1.46 (0.10) 1.16 (0.07) 0.94 (0.06) 0.87 (0.06) 0.85 (0.05) 0.82 (0.05)
KRR-$\widehat{\mathbf{Z}}$ 2.83 (0.15) 1.56 (0.13) 1.14 (0.06) 0.93 (0.06) 0.86 (0.05) 0.84 (0.05) 0.82 (0.05)
KRR-$\mathbf{Z}$ 0.74 (0.04) 0.75 (0.04) 0.74 (0.04) 0.74 (0.04) 0.74 (0.04) 0.74 (0.04) 0.74 (0.04)
KRR-$\mathbf{X}$ 3.79 (0.16) 3.74 (0.16) 3.47 (0.14) 2.53 (0.11) 2.11 (0.10) 1.88 (0.08) 1.39 (0.07)
LR-$\mathbf{Z}$ 3.62 (0.14) 3.59 (0.15) 3.61 (0.14) 3.59 (0.14) 3.58 (0.13) 3.61 (0.15) 3.61 (0.15)
As shown in the first two panels of Table 3, the proposed KRR-$\widehat{\mathbf{Z}}^{\prime}$ outperforms KRR-$\mathbf{X}$ and LR-$\mathbf{Z}$. Comparing to the oracle approach $\mathbf{Z}, KRR$-$\widehat{\mathbf{Z}}^{\prime}$ has increasingly closer performance as $n$ or $p$ grows. This is in line with our risk bound in Corollary 5 where the extra prediction error of KRR-$\widehat{\mathbf{Z}}^{\prime}$ relative to KRR-$\mathbf{Z}$ is quantified by $\mathcal{O}_{\mathbb{P}}(1/p+1/n)$, which vanishes with increasing $n$ and $p$. Comparing to KRR-$\widehat{\mathbf{Z}}$ that does not use auxiliary data, the benefit of using auxiliary data in KRR-$\widehat{\mathbf{Z}}^{\prime}$ is evident for small sample sizes. As $n$ increases, the prediction error of KRR-$\mathbf{X}$ does not decrease as fast as those of KRR-$\widehat{\mathbf{Z}}^{\prime}$ and KRR-$\mathbf{Z}$, indicating a slower rate of convergence of the former. This is due to that KRR-$\mathbf{X}$ does not exploit the low-dimensional structure. As $p$ increases, we also note that the performance of KRR-$\mathbf{X}$ slightly increases. One explanation could be that more useful information for predicting $Y$ can be found from $\mathbf{X}$ as $p$ increases. Since KRR-$\mathbf{Z}$ and LR-$\mathbf{Z}$ use $\mathbf{Z}$ instead of $\mathbf{X}$, their prediction errors remain the same as $p$ increases.
Regarding the effect of SNR, a higher SNR implies that $\mathbf{X}$ contains more information about $\mathbf{Z}$, which should lead to better prediction of $\mathbf{Z}$ and consequently of $Y$. As seen in the third panel of Table 3, the predictive performance of $\widehat{\mathbf{Z}}^{\prime}, KRR-$$\widehat{\mathbf{Z}}$, and KRR-$\mathbf{X}$ indeed improves as the SNR increases, whereas the prediction errors of KRR-$\mathbf{Z}$ and LS-$\mathbf{Z}$ remain unchanged. For large values of SNR, KRR-$\widehat{\mathbf{Z}}^{\prime}$ performs very well, approaching to KRR-$\mathbf{Z}$, whereas KRR-$\mathbf{X}$ still performs poorly, due to the fact that it does not fully exploit the low-dimensional structure. We also observe from the comparison between KRR-$\widehat{\mathbf{Z}}^{\prime}$ and KRR-$\widehat{\mathbf{Z}}$ that the benefit of using auxiliary data becomes more visible for small SNRs.
Finally, comparing to LS-$\mathbf{Z}$ in all settings, KRR-$\widehat{\mathbf{Z}}^{\prime}$ has substantial predictive advantage, implying the benefit of using KRR to capture the non-linear relationship between the response and the latent factor.
7 Real data analysis
In this section, we present real data analysis on two datasets based on pretrained LLM embeddings, one tailored to a regression problem and the other to a classification problem.
7.1 Analysis on Letterboxd Film Dataset
In this section, we investigate the predictive performance of KRR using pre-trained embedding vectors from movie reviews in the Letterboxd Film Dataset (Canseven, 2025). The dataset contains information on 847,209 films from the Letterboxd platform, each with multiple textual reviews and an average user rating on a 1–5 scale. Our goal is to train a predictive model that can accurately predict the rating score from the reviews of a given film.
Unlike the classical regression setting, this task involves textual inputs that must be converted to numerical embedding vectors before applying KRR. We follow the procedure in Example 3 of the Introduction, first applying a pretrained language model to transform the textual review content into embedding vectors in $\mathbb{R}^{d}$. To this end, we choose the General Text Embeddings (GTE) model Li et al. (2023), which is trained based on the BERT framework Devlin et al. (2019). By leveraging multi-stage contrastive learning, the GTE model efficiently captures sentence-level similarity in order to generate text embedding representations, rendering it broadly applicable to various natural language tasks. The GTE model provides three different sizes of models: thenlper/gte-small, thenlper/gte-base, and thenlper/gte-large, producing embedding vectors with $d=384,768$, and $1024$, respectively. In addition to generating text embeddings, we also investigate whether further exploiting the low dimensional structure of the embedding vectors by applying PCA to reduce their dimensions offers additional benefits. Combining these two steps yields the final predicted inputs (see, Figure 2, for an illustration), which are used in the downstream regression task, with the rating score of each movie review as the response.
Great movie, but the story felt flat. -0.02489 -0.05562 $\vdots$ 0.02846 0.8621 -0.0542 0.0841 EmbeddingPCA
Figure 2: An illustration of the process generating predicted feature inputs with $r=3$
We consider three methods to predict the rating scores using the predicted inputs: (1) KRR with a Gaussian kernel, (2) KRR with a linear kernel, and (3) ordinary least squares (OLS). We study the effect of the embedding dimension $d$ by selecting different GTE models. We also vary the retained low-dimension $r$ in (44), ranging from 1 to $d$. We randomly select from the whole dataset 13864 sub-samples and for each setting, we randomly select $2000$ samples for training and another $2000$ samples for testing. The remaining samples are used in PCA to construct $\widehat{B}$ as in (45). By repeating 50 times in each setting, we report the averaged mean squared errors (MSE) on the test data in Figure 3.
Figure 3: Averaged MSE of different methods with varying $r$ and GTE models
There are several interesting observations that can be made. First, we see that for all methods, increasing the dimension of embedding vectors leads to better predictive performance. This is in line with our intuition that pre-trained embedding vectors with larger dimension tend to encode more useful information for prediction. Second, for a given $d$, linear KRR with $r=d$ has the smallest MSE, suggesting a linear relationship between the embedding vectors and the ratings. It also indicates that such a linear hyperplane in the embedding space cannot be fully explained by its leading principal components. The other linear predictor, OLS, achieves its best performance at an intermediate $r$, striking a balance between variance and bias, as shown in Corollary 5 of Bing et al. (2021). In contrast, linear KRR exhibits a clear advantage for relatively large $r$ because the added regularization helps prevent overfitting, in line with Theorem 1 and Corollary 2. Third, the Gaussian KRR exhibits interesting predictive performance as $r$ varies. For relatively small $r$, KRR with Gaussian kernel in general outperforms the other methods. This is expected as linear methods are insufficient to characterize the dependence of responses on only a few features, leading to under-fitting, whereas KRR with Gaussian kernel is more flexible in capturing the non-linear relationship. However, unlike KRR with a linear kernel, whose performance consistently improves as $r$ increases, KRR with a Gaussian kernel shows a slight decline starting at $r=32$. One explanation is that as $r$ grows beyond this point, the model begins to incorporate principal components less relevant to the signal (also evident from the plateau of the MSE for the linear kernel), leading to larger errors in predicting the features. These errors are amplified by the nonlinear Gaussian kernel, whereas the linear kernel is less sensitive to them. As our theory suggests, the impact of feature-prediction errors depends on the kernel, though a more refined analysis is needed to fully characterize this dependence, which we leave to future work. Finally, when $r$ approaches $d$, the last principal components contribute additional signal, and the Gaussian kernel achieves performance comparable to linear KRR. Our overall recommendation is to use linear KRR when the goal is to achieve the best predictive performance with the full embedding, and to use Gaussian KRR when one also seeks to obtain a small set of extracted features for interpretability.
7.2 Analysis on Rotten Tomatoes Movie Review Dataset
In this section, we validate the kernel learning method with predicted feature inputs on Rotten Tomatoes Movie Review Dataset (Pang and Lee, 2005), which contains 5331 positive and 5331 negative reviews. This benchmark dataset has been widely employed in sentiment categorization research. For this binary classification task, the loss function in (48) is specified as the hinge loss, corresponding to the kernel support machine vector (KSVM) method.
Similar to the treatment in the text regression analysis in Section 7.1, we first use pretrained GTE models to generate text embedding vectors, and then apply PCA as in (44) to obtain the final predicted feature inputs. For comparison, we consider three classifiers for sentiment categorization using predicted inputs: (1) KSVM with a Gaussian kernel; (2) KSVM with a linear kernel; (3) logistic regression. The selection of the bandwidth for Gaussian kernel and the regularization parameter $\lambda$ in (48) follows the same procedure as in Section 7.1, We vary the dimension of the text embedding vectors by selecting different GTE models from {thenlper/gte-small, thenlper/gte-base, thenlper/gte-large}, and the retained low-dimension $r$ in PCA from 1 to the full embedding dimension $d$. For each setting, we randomly select from the overall dataset 1000 samples for training and another 1000 samples for testing, and the remaining data is used to perform PCA. By repeating 50 times in each setting, the averaged accuracy rates are reported in Figure 4.
Figure 4: Averaged accuracy rates of different methods with varying $r$ and GTE models
Some findings from Figure 4 are in order. First, increasing the dimension of embedding vectors generally leads to an increase in classification accuracy. This is expected as the embedding vectors with larger dimension tend to encode more useful information for classification. Second, the performance between KSVM classifiers with Gaussian and linear kernels are nearly identical across all settings. Moreover, their accuracy rate increases substantially as $r$ changes from $1$ to $2$, after which it rapidly stabilizes to reach their best performance. This indicates that the relationship between embedding vectors and the labels is linear and can be fully explained by the first few leading principal components. Third, for large $r$, the logistic classifier suffers from overfitting, leading to unsatisfactory performance in classification. This is consistent with the behavior of OLS observed from text regression analysis in Section 7.1. Overall, our experiments suggest that, to avoid tuning the optimal $r$, KSVM with the full embedding—whether using a linear or Gaussian kernel—is the preferred choice for the text classification task, although using $r=2$ yields only slightly worse accuracy while remaining interpretable.
@@ 7.2 Analysis on Rotten Tomatoes Movie Review Dataset
The main proofs are collected in Appendix A and presented on a section-by-section basis. Auxiliary lemmas are stated in Appendix B.
Appendix A Main Proofs